# Feature Importance-Aware Deep Joint Source-Channel Coding for Computationally Efficient and Adjustable Image Transmission

### v2

@@ Abstract
Recent advances in deep learning-based joint source–channel coding (deepJSCC) have substantially improved communication performance, but their high computational cost hinders practical deployment. Moreover, certain applications require the ability to dynamically adapt computational complexity. To address these issues, we propose a Feature Importance-Aware deepJSCC (FAJSCC) model for image transmission that is both computationally efficient and adjustable. FAJSCC employs axis-dimension specialized computation, which performs efficient operations individually for each spatial and channel axis, significantly reducing computational cost while representing features effectively. It further incorporates selective deformable self-attention, which applies self-attention only to selected and adaptively adjusted regions, leveraging the importance and relations of input features to efficiently capture complex feature correlations. Another key feature of FAJSCC is that the number of selected important areas can be controlled separately by the encoder and the decoder, depending on the available computational budget. It makes FAJSCC the first deepJSCC architecture to allow independent adjustment of encoder and decoder complexity within a single trained model. Experimental results show that FAJSCC achieves superior image transmission performance under various channel conditions while requiring less computational complexity than recent state-of-the-art models. Furthermore, experiments independently varying the encoder and decoder’s computational resources reveal, for the first time in the deepJSCC literature, that understanding the meaning of noisy features in the decoder demands the greatest computational cost. The code is publicly available at github.com/hansung-choi/FAJSCCv2.
Index Terms:
joint source-channel coding, feature importance, image transmission, computational complexity
@@ I Introduction
Conventional digital communication systems rely on separation-based source and channel coding according to Shannon’s source-channel coding separation theorem [1], where source coding compresses data to reduce its size, and channel coding adds redundancy to make data robust against channel noise without losing optimality. Based on this modularity, source coding (e.g., JPEG, JPEG2000, PNG, BPG, and AVIF) and channel coding methods (e.g., Turbo, LDPC, and Polar codes) have been developed independently for several decades. However, it is known that the joint source-channel coding (JSCC) design strictly outperforms the separation-based design in several scenarios, such as finite block length [2] and sending correlated data over multi-user settings [3]. Early analytic research on the JSCC was limited to simple data distributions, like Gaussian cases, due to the tractability, and hence could not deal with real data, such as images [4, 5].
@@
One of the breakthroughs in advancing the JSCC beyond its early results is the use of deep learning, known as deepJSCC, which has shown significant performance improvement for image transmission [6]. Since the introduction of deepJSCC, extensive efforts have been made, leading to great success in terms of communication performance. For instance, deepJSCC models have shown better data reconstruction performance compared to separation-based systems for simple point-to-point additive white Gaussian noise (AWGN) channels [6, 7, 8, 9, 10], various fading channels  [11, 12, 13, 14, 15, 16], and several bandwidth settings [17, 18, 19] as well as diverse network configurations such as MIMO [20] and relay networks [21]. Such deepJSCC models for image transmission are also extended for other data modalities such as speech [22] and text [23]. Recent semantic communication can be also viewed as a variant of deepJSCC designed for downstream tasks, such as image retrieval [24], segmentation [25], and multiple tasks [26, 27].
@@
Despite its success in communication performance, existing deepJSCC research is increasingly suffering from growing computational complexity incurred by large neural networks, along with associated problems like high power consumption, hardware costs, and latency, which prohibit its practical deployment. To reduce such computational burden, one common approach is to apply model compression techniques. For instance, pruning [28, 29, 30] and low-rank decomposition [31] compress trained models, successfully reducing the number of parameters (i.e., computational load) at the expense of performance degradation and extra training efforts.
@@
Another important function that is often required in applications is the ability to dynamically control computational complexity within a single model. For instance, in surveillance applications, it is essential to transmit images of large surveillance areas at low computational complexity (i.e., low resolution) to ensure minimal delay and power consumption for long-term monitoring, while transmitting at high computational complexity (i.e., high resolution) when detailed analysis is needed. To address the varying computational needs at the application level, it is essential to dynamically manage the tradeoff between computational complexity and data communication accuracy. However, to the best of our knowledge, this issue is largely unstudied for image transmission in existing literature.
@@ III-B Selective Deformable Self-Attention
Here, $(u,v)$ denotes the deformed sampling location predicted from the original coordinate $(i,j)$ and the predicted offset $\mathbf{z}_{{\textrm{offset}}}\in\mathbb{R}^{H_{{\textrm{in}}}\times W_{{\textrm{in}}}\times 2}$ from the input feature relations. The four neighboring integer grid points $i_{0},i_{1},j_{0},j_{1}$ represent the surrounding pixels (or features) that are closest to $(u,v)$. The interpolation weights $\alpha$ and $\beta$ correspond to the fractional distances between $(u,v)$ and the integer coordinates along the height and width axes, respectively. The equation for $z_{{\textrm{out}}}^{c}(i,j)$ performs bilinear interpolation: When $(u,v)$ coincides with an integer grid point, it directly samples the corresponding feature value, whereas when $(u,v)$ lies between grid points, it computes a weighted combination of the four surrounding features, thereby interpolating the feature map at the non-integer position. This formulation ensures that both cases (direct alignment with existing feature coordinates and interpolation at in-between positions) are handled within a single equation, allowing subsequent self-attention to adaptively operate on adjusted regions instead of being limited to fixed grids. This effect is illustrated in Figure 4.111Note that the figure (and subsequent figures as well) illustrates the deformation in the image domain to ease understanding, but the actual deformation operates in the feature domain. The thick black borders in this figure indicate the features selected for self-attention. Applying self-attention within fixed-window areas may overlook feature correlations beyond them. In contrast, feature deformation enables self-attention to operate on adaptively adjusted regions, thereby capturing correlations that go beyond the fixed windows. We call this method deformable (window) self-attention.
@@
Deformable self-attention effectively enhances the capability of standard self-attention. Building on this, we propose to applying deformable self-attention selectively to improve computational efficiency. Previous studies in neural source coding have shown that not all features carry the same amount of information: Some features are more informative (e.g., main objects), while others are less informative (e.g., background regions) [42]. Moreover, it has been demonstrated that simple images do not require high computational cost to achieve high performance in super-resolution tasks [43, 44]. Motivated by these observations, we apply deformable self-attention only to features in the most important windows, identified based on their contribution to image transmission performance. By restricting self-attention to these windows, the computational cost is significantly reduced. This selective enhancement strategy greatly reduces computational cost compared with applying deformable self-attention across all windows. A portion of the saved computation is then reinvested to increase intermediate feature sizes, leading to richer representations. Although this reinvestment slightly raises the complexity, the increase is much smaller than the computational savings, so the overall cost remains significantly lower than applying self-attention to features in all windows. As a result, FAJSCC achieves higher transmission performance by enhancing important feature processing while still maintaining reduced overall complexity.
@@ III-C Attention Family Tree
The proposed axis-dimension specialized operation and the selective deformable self-attention are both highly efficient and impose minimal computational overhead. However, we observe that naïvely computing the attention features, offsets, and window importance from scratch negates the efficiency gains of our methods; thus, we propose a new attention family tree that jointly extracts attention features and obtains offset and feature importance at once by eliminating overlapping calculations based on relations among output features to minimize computational resource overhead.
@@ III-D Training and Loss Function Design
The hard sampling result $\mathbf{z}_{{\textrm{hard}}}$ follows the exact true distribution with $\mathbf{\pi}^{1},\mathbf{\pi}^{2}$ probabilities, but this prevents backpropagation. In contrast, the soft sampling result $\mathbf{z}_{{\textrm{soft}}}$ enables backpropagation, but it deviates to some extent from the true distribution, except when $\tau=0$, i.e., $\mathbf{z}_{{\textrm{soft}}}=\mathbf{z}_{{\textrm{hard}}}$. To leverage the advantages of $\mathbf{z}_{{\textrm{soft}}},\mathbf{z}_{{\textrm{hard}}}$ both, we use the straight-through important window sampling result $\mathbf{z}_{{\textrm{ST}}}$ as follows:
@@ IV-A Experimental Setting
Baselines and Architectures: We implement two versions of FAJSCC: The first is as illustrated in Figure 2 with details in Section III. The other is without the selective deformable self-attention proposed in Section III-B, i.e., unused offset and window importance heads in the attention family tree, and the selective deformable self-attention module in the main computational path are removed. We call the latter version light-attention deepJSCC (LAJSCC). LAJSCC uses only the axis-dimension specialized computation proposed in Section III-A to verify its computational efficiency. The other deepJSCC baselines are ranging from the original deepJSCC [6] to the recent SwinJSCC [8] and LICRFJSCC [32]. To differentiate between these baselines, we refer to the initial deepJSCC [6] based on the convolution blocks and the deepJSCC based on residual convolution blocks [7] as ConvJSCC and ResJSCC, respectively. The detailed implementation code is available at github.com/hansung-choi/FAJSCCv2.
@@ IV-B Main Result
TABLE I: Comparison for GFLOPs and memory usage of deepJSCCs under AWGN channel, SNR=10dB10\ $\mathrm${dB}, and CPP=$1/16$.
@@
ConvJSCC [6] 22.50 4.48 0.64
ResJSCC [7] 25.51 5.73 0.77
SwinJSCC [8] 26.46 20.02 71.26
LICRFJSCC [32] 25.61 9.02 3.84
LAJSCC (ours) 27.36 11.35 5.27
FAJSCC (ours) 27.92 17.58 13.56
@@
Architecture Efficiency: Table I shows PSNR performance in terms of computational cost (GFLOPs) and memory usage (MB) under AWGN channel, SNR=1010dB, and CPP=$1/16$. Here, GFLOPs are measured when transmitting $512\times 768$ resolution image. In this table, our LAJSCC and FAJSCC achieve the highest PSNR while using smaller computational resources and memory than SwinJSCC [8].
PSNR and SSIM Results: Figures 6 and 7 show the PSNR and SSIM performance results across various communication environments. Our LAJSCC and FAJSCC outperform the previous models for different performance metrics under various communication environments. This performance gain primarily comes from our dimension-specialized computation approaches: (1) Using only axis-dimension specialized operations (LAJSCC) shows higher performance than the recent SOTA SwinJSCC in most communication environments, while requiring computational resources about half and $14$ times smaller memory. This is a notable achievement, as even the recent SOTA lightweight model LICRFJSCC [32] reduces computational resources by shrinking feature sizes when applying self-attention, which inherently limits its performance, making it impossible to reach that of SwinJSCC. (2) Upon the axis-dimension specialized operations, our proposed selective deformable self-attention (FAJSCC) even more enhances the image transmission performance by utilizing deformation and feature importance. These advantages lead FAJSCC to outperform SwinJSCC while using smaller computational and memory resources.

### v3

@@ Abstract
Recent advances in deep learning-based joint source–channel coding (deepJSCC) have substantially improved communication performance, but their high computational cost hinders practical deployment. Moreover, certain applications require the ability to dynamically adapt computational complexity. To address these issues, we propose a Feature Importance-Aware deepJSCC (FAJSCC) model for image transmission that is both computationally efficient and adjustable. FAJSCC employs axis-dimension specialized computation, which performs efficient operations individually for each spatial and channel axis, significantly reducing computational cost while representing features effectively. It further incorporates selective deformable self-attention, which applies self-attention only to selected and adaptively adjusted features, leveraging the importance and relations of input features to efficiently capture complex feature correlations. Another key feature of FAJSCC is that the number of selected important areas can be controlled separately by the encoder and the decoder, depending on the available computational budget. It makes FAJSCC the first deepJSCC architecture to allow independent adjustment of encoder and decoder complexity within a single trained model. Experimental results show that FAJSCC achieves superior image transmission performance under various channel conditions while requiring less computational complexity than recent state-of-the-art models. Furthermore, experiments independently varying the encoder and decoder’s computational resources reveal, for the first time in the deepJSCC literature, that understanding the meaning of noisy features in the decoder demands the greatest computational cost. The code is publicly available at github.com/hansung-choi/FAJSCCv2.
@@ I Introduction
Advances in high-performance embedded hardware and next-generation wireless communication technologies such as 6G have enabled the widespread deployment of IoT devices across diverse application domains, including smart healthcare, smart manufacturing, and smart cities. In particular, vision IoT devices, such as wireless surveillance cameras and UAVs, provide important visual information for security and safety monitoring, including real-time intrusion detection and inspection of hazardous areas. However, the rapid growth of IoT devices has increased pressure on available bandwidth. Especially, vision IoT devices require substantial bandwidth to transmit visual data, which are inherently more complex and contain a much larger amount of information, thereby requiring significantly more bits than other modalities such as text and speech. Therefore, the recent communication systems need high-quality vision data transmission methods under limited bandwidth constraints.
@@
Although conventional separation-based digital communication systems have undergone rapid improvements in both source coding (e.g., JPEG, JPEG2000, PNG, BPG, and AVIF) and channel coding (e.g., Turbo, LDPC, and polar codes) over several decades, further manual improvements to these rule-based methods have become increasingly difficult in satisfying recent communication requirements. In other side, it is known that the joint source-channel coding (JSCC) design strictly outperforms the separation-based design in several scenarios, such as finite block length [1] and sending correlated data over multi-user settings [2]. Early analytic research on the JSCC was limited to simple data distributions, like Gaussian cases, due to the tractability, and hence could not deal with real data, such as images [3, 4].
@@
One of the breakthroughs in advancing the JSCC beyond its early results is the use of deep learning, known as deepJSCC, which has shown significant performance improvement for image transmission [5] and become close to satisfying recent communication requirements. Since the introduction of deepJSCC, extensive efforts have been made, leading to great success in terms of communication performance. For instance, deepJSCC models have shown better data reconstruction performance compared to separation-based systems for simple point-to-point additive white Gaussian noise (AWGN) channels [5, 6, 7, 8, 9], various fading channels  [10, 11, 12, 13, 14, 15], and several bandwidth settings [16, 17, 18] as well as diverse network configurations such as MIMO [19] and relay networks [20]. Recent vision-related semantic communication can also be viewed as a variant of deepJSCC designed for downstream tasks, such as image retrieval [21], segmentation [22], and multiple tasks [23, 24].
@@
Despite its success in communication performance, existing deepJSCC research is increasingly suffering from growing computational complexity incurred by large neural networks, along with associated problems like high power consumption, hardware costs, and latency, which prohibit its practical deployment, especially for IoT devices. To reduce such computational burden, one common approach is to apply model compression techniques. For instance, pruning [25, 26, 27] and low-rank decomposition [28] compress trained models, successfully reducing the number of parameters (i.e., computational load) at the expense of performance degradation and extra training efforts.
@@
Another important function that is often required in applications is the ability to dynamically control computational complexity within a single model. For instance, in vision monitoring applications of IoT devices, it is essential to transmit images of large areas at low computational complexity (i.e., low resolution) to ensure minimal delay and power consumption for long-term monitoring, while transmitting at high computational complexity (i.e., high resolution) when detailed analysis is needed. To address the varying computational needs at the application levels, it is essential to dynamically manage the tradeoff between computational complexity and data communication accuracy. However, to the best of our knowledge, this issue is largely unstudied for image transmission in existing deepJSCC literature.
@@ III-B Selective Deformable Self-Attention
Here, $(u,v)$ denotes the deformed sampling location predicted from the original coordinate $(i,j)$ and the predicted offset $\mathbf{z}_{{\textrm{offset}}}\in\mathbb{R}^{H_{{\textrm{in}}}\times W_{{\textrm{in}}}\times 2}$ from the input feature relations. The four neighboring integer grid points $i_{0},i_{1},j_{0},j_{1}$ represent the surrounding pixels (or features) that are closest to $(u,v)$. The interpolation weights $\alpha$ and $\beta$ correspond to the fractional distances between $(u,v)$ and the integer coordinates along the height and width axes, respectively. [^1] The interpolation weights $\alpha$ and $\beta$ enable gradient flow into $\mathbf{z}_{{\textrm{offset}}}$. The equation for $z_{{\textrm{out}}}^{c}(i,j)$ performs bilinear interpolation: When $(u,v)$ coincides with an integer grid point, it directly samples the corresponding feature value, whereas when $(u,v)$ lies between grid points, it computes a weighted combination of the four surrounding features, thereby interpolating the feature map at the non-integer position. This formulation ensures that both cases (direct alignment with existing feature coordinates and interpolation at in-between positions) are handled within a single equation, allowing subsequent self-attention to adaptively operate on adjusted regions instead of being limited to fixed grids. This effect is illustrated in Figure 4.222Note that the figure (and subsequent figures as well) illustrates the deformation in the image domain to ease understanding, but the actual deformation operates in the feature domain. The thick black borders in this figure indicate the features selected for self-attention. Applying self-attention within fixed-window areas may overlook feature correlations beyond them. In contrast, feature deformation enables self-attention to operate on adaptively adjusted regions, thereby capturing correlations that go beyond the fixed windows. We call this method deformable (window) self-attention.
@@
Deformable self-attention effectively enhances the capability of standard self-attention. Building on this, we propose to applying deformable self-attention selectively to improve computational efficiency. Previous studies in neural source coding have shown that not all features carry the same amount of information: Some features are more informative (e.g., main objects), while others are less informative (e.g., background regions) [39]. Moreover, it has been demonstrated that simple images do not require high computational cost to achieve high performance in super-resolution and image compression tasks [40, 47, 41]. Motivated by these observations, we apply deformable self-attention only to features in the most important windows, identified based on their contribution to image transmission performance. By restricting self-attention to these windows, the computational cost is significantly reduced. This selective enhancement strategy greatly reduces computational cost compared with applying deformable self-attention across all windows. A portion of the saved computation is then reinvested to increase intermediate feature sizes, leading to richer representations. Although this reinvestment slightly raises the complexity, the increase is much smaller than the computational savings, so the overall cost remains significantly lower than applying self-attention to features in all windows. As a result, FAJSCC achieves higher transmission performance by enhancing important feature processing while still maintaining reduced overall complexity.
@@ III-C Attention Family Tree
The proposed axis-dimension specialized operation and the selective deformable self-attention are both highly efficient and impose minimal computational overhead. However, we observe that naïvely computing the attention features, offsets, and window importance from scratch negates the efficiency gains of our methods; thus, we propose a new attention family tree that jointly extracts attention features and obtains offset and feature importance at once by eliminating overlapping calculations based on relations among output features to minimize computational resource overhead. The attention family tree structure is shown in Figure 3.
@@ III-D Training and Loss Function Design
The hard sampling result $\mathbf{z}_{{\textrm{hard}}}$ follows the exact true distribution with $\mathbf{\pi}^{1},\mathbf{\pi}^{2}$ probabilities, but this prevents backpropagation. In contrast, the soft sampling result $\mathbf{z}_{{\textrm{soft}}}$ enables backpropagation, but it deviates to some extent from the true distribution, except when $\tau=0$, i.e., $\mathbf{z}_{{\textrm{soft}}}=\mathbf{z}_{{\textrm{hard}}}$. To leverage the advantages of $\mathbf{z}_{{\textrm{soft}}},\mathbf{z}_{{\textrm{hard}}}$ both, we use the straight-through important window sampling result $\mathbf{z}_{{\textrm{ST}}}$ as follows: [^4] Straight-through is a technique for approximating gradients of non-differentiable functions to enable backward propagation.
@@ IV-A Experimental Setting
Baselines and Architectures: We implement two versions of FAJSCC: The first is as illustrated in Figure 2 with details in Section III. The other is without the selective deformable self-attention proposed in Section III-B, i.e., unused offset and window importance heads in the attention family tree, and the selective deformable self-attention module in the main computational path are removed. We call the latter version light-attention deepJSCC (LAJSCC). LAJSCC uses only the axis-dimension specialized computation proposed in Section III-A to verify its computational efficiency. The other deepJSCC baselines range from the original deepJSCC [5] to the recent SwinJSCC [7] and LICRFJSCC [29].
To differentiate between these baselines, we refer to the initial deepJSCC [5] based on the convolution blocks and the deepJSCC based on residual convolution blocks [6] as ConvJSCC and ResJSCC, respectively. In contrast to SwinJSCC [7] that applies self-attention to all features, LICRFJSCC [29] applies self-attention only to a fixed subset of feature channels. Unless explicitly specified otherwise, the GFLOPs and memory usages of implemented models follow Table I. The scaling settings for ConvJSCC and ResJSCC follow those in [5], while the scaling configuration for SwinJSCC is determined via an exhaustive search to achieve the best possible performance under comparable GFLOPs constraints. LICRFJSCC, LAJSCC, and FAJSCC scale settings are based on SwinJSCC. The detailed implementation code is available at github.com/hansung-choi/FAJSCCv2.
@@
Figure 6: Efficiency comparison for various model sizes with respect to computational burden (GFLOPs) and memory usage.
@@ IV-B Main Result
TABLE I: Comparison for GFLOPs and memory usage of deepJSCCs under AWGN channel, SNR=10dB10\ $\mathrm${dB}, and CPP=$1/32$.
@@
ConvJSCC [5] 21.09 4.48 0.64
ResJSCC [6] 24.05 5.73 0.77
SwinJSCC [7] 25.90 20.02 71.26
LICRFJSCC [29] 24.52 9.02 3.84
LAJSCC (ours) 26.55 11.35 5.27
FAJSCC (ours) 26.86 17.58 13.56
@@
Architecture Efficiency: Table I shows PSNR performance in terms of computational cost (GFLOPs) and memory usage (MB) under AWGN channel, SNR=1010dB, and CPP=$1/32$. Here, GFLOPs are measured when transmitting $512\times 768$ resolution image. In this table, our LAJSCC and FAJSCC achieve the highest PSNR while using smaller computational resources and memory than SwinJSCC [7].
To fully verify the performance gain of our model, we provide efficiency comparison for various model sizes in Figure 6. To match the computational resource usage for comparison, we implement smaller and larger versions of SwinJSCC, LAJSCC, and FAJSCC by decreasing and increasing feature sizes by about $50\%$. In the case of ConvJSCC and ResJSCC, we increase the feature size by factors of $2$ and $4$ to construct larger models.
Figure 6 shows the tradeoff between computational block complexity (GFLOPs and memory usage) and PSNR performance for various JSCC schemes. Most existing works, such as ConvJSCC [5], ResJSCC [6], and SwinJSCC [7], primarily aim to improve PSNR. These performance gains are achieved by employing heavier computational blocks, which consequently incur substantially higher GFLOPs and memory requirements. In particular, SwinJSCC demands approximately an order of magnitude more memory than ConvJSCC and ResJSCC, and imposes a significantly higher minimum computational burden for the same feature scales due to its reliance on heavy multi-head self-attention mechanisms. This increased requirement gives a large burden to practical deployment, especially for IoT devices. To address this issue, recent works decrease computational burden at the expense of performance degradation. For example, LICRFJSCC [29] applies advanced self-attention only to a fixed subset of feature channels without considering feature importance. Although this method reduces the minimum level of computational burden, it also compromises performance by shrinking the feature representation space.
Different from such previous directions, our method achieves performance improvement while reducing computational burden, i.e., going to the upper left in the left plot of Figure 6. Our FAJSCC is located in the far upper-left corner by applying computationally expensive self-attention only to the important features. Moreover, by disabling the selectively deformable self-attention block, the FAJSCC model becomes LAJSCC. In this figure, FAJSCC’s PSNR vs. GFLOPs is similar to that of the LAJSCC. This verifies that FAJSCC can dynamically control computational resources without compromising the performance efficiency gain in PSNR vs. GFLOPs achieved by LAJSCC.
PSNR and SSIM Results: Figures 7 and 8 show the PSNR and SSIM performance results across various communication environments. Our LAJSCC and FAJSCC outperform the previous models for different performance metrics under various communication environments. This performance gain primarily comes from our dimension-specialized computation approaches: (1) Using only axis-dimension specialized operations (LAJSCC) shows higher performance than the recent SOTA SwinJSCC in most communication environments, while requiring computational resources about half and $14$ times smaller memory. This is a notable achievement, as even the recent SOTA lightweight model LICRFJSCC [29] reduces computational resources by shrinking feature sizes when applying self-attention, which inherently limits its performance, making it impossible to reach that of SwinJSCC. (2) Upon the axis-dimension specialized operations, our proposed selective deformable self-attention (FAJSCC) even more enhances the image transmission performance by utilizing deformation and feature importance. These advantages lead FAJSCC to outperform SwinJSCC while using smaller computational and memory resources.